Prove that 2 to the power n+1 > (n + 2) · sin(n) for all positive integers n. 


1
Expert's answer
2021-08-16T16:53:02-0400

Consider the function "f(x)=2^{x+1}-(x+2)."

"f'(x)=2^{x+1}\ln2-1\geq2^{1+1}\ln2-1=\ln(\dfrac{4}{e})>0, x\geq1"

Then the function "f(x)" increases for "x\geq1."

"f(1)=2^{1+1}-(1+2)=1>0"

Hence


"2^{n+1}>(n+2), n\geq 1""1\geq\sin n, n>1"

Then


"2^{n+1}>n+2\geq(n+2)\cdot\sin n, n\geq1"


Therefore "2^{n+1}>(n+2)\cdot\sin n," for all positive integers "n."